OpenStudy (anonymous):

Ken has decided to raise rabbits but has been warned that the number of rabbits will double every month. ken started out with 30 rabbits and the function y=30x 2^x models the number of rabits he will have after x months. What graph represents this function ?

OpenStudy (anonymous):

Those are the graphs

OpenStudy (anonymous):

Easiest way to find the correct graph is to plug in a few x values and see which graph they fit. Remember that raising 2 to the power of x means that as x goes up you are doubling an even bigger number than the previous so it tends to go up quickly also, it is not constant since doubling 30 is much different than doubling 120.

OpenStudy (anonymous):

is it the 1st one?

OpenStudy (anonymous):

@Dmadaras

OpenStudy (anonymous):

Yes. That is the closest approx. graph since 30*2^4 = 30 * 16 = 580 when x= 4, and y = 1160 when x = 5 .

OpenStudy (anonymous):

Thank you very much.